# Reconfigurable Intelligent Surface-Assisted Localization in OFDM Systems with Carrier Frequency Offset and Phase Noise

### v3

@@ Abstract
Reconfigurable intelligent surface (RIS)-assisted communication systems have been extensively studied for providing high-precision location services. However, most studies have overlooked the impact of carrier frequency offset (CFO) and phase noise (PN) resulting from hardware impairments on localization. This paper presents a novel, alternating optimization (AO)-based algorithm to jointly estimate the CFO, PN, and user position, where, provided the user position, closed-form expressions for the CFO and PN are derived per iteration, significantly reducing the complexity and enhancing the stability of the algorithm. Another important aspect is a new RIS phase shift optimization algorithm developed to minimize the analytical lower bound of localization accuracy, hence benefiting localization. The semidefinite relaxation method and Schur complement are utilized to convexify this challenging non-convex optimization problem to a semidefinite program. Simulations demonstrate the effectiveness of the proposed algorithms, with the localization accuracy enhanced by two orders of magnitude. The localization accuracy of the proposed algorithm is close to the analytical lower bound, with a root mean square error of lower than $\rm 10^{-2}\>m$.
@@ I-A Motivation and Challenges
With the development of industrial Internet of Things, there is an increasing demand for centimeter-level localization accuracy in the sixth-generation (6G) mobile networks [1, 2, 3]. While global navigation satellite systems (GNSSs) have been widely applied to outdoor scenarios [4], they undergo insufficient localization accuracy and signal instability arising from obstacles [5, 6]. In recent years, communication systems have been considered to assist in localization, e.g., integrated sensing and communication (ISAC), aiming to achieve more accurate and continuous localization alongside GNSS signals.
Although communication-assisted localization systems have been extensively studied, few studies have considered the adverse effects of hardware impairments, such as carrier frequency offset (CFO) and phase noise (PN), on localization accuracy. CFO is a constant parameter due to the frequency offset between the transmitter and the receiver. Doppler shift can also lead to CFO. PN is a random process resulting from the fluctuation of the oscillators at the transmitters and receivers [7]. CFO and PN can result in signal constellation rotations in single carrier communication systems, or inter-carrier interference (ICI) in broadband communication systems, for example, orthogonal frequency division multiplexing (OFDM) systems. The overall communication performance and the localization accuracy can degrade significantly, due to the CFO and PN.
Another challenge arises from the increasingly likely blockages of signals, e.g., millimeter wave (mmWave) signals with the quasi-optical property, in many dense urban areas [8]. Reconfigurable intelligent surface (RIS) [9, 10] has been considered in communication systems to deal with the problem of signal blockage. Localization can be implemented with the help of virtual line-of-sight (LoS) paths provided by RISs even when the LoS paths are blocked, and can be improved via passive beamforming. In this sense, RIS-assisted user localization is of practical interest, and could be particularly challenging with the further consideration of the CFO and PN. However, this has not been studied in the existing literature.
@@ I-B Related works
Communication systems have been widely employed to assist in high-precision user localization, typically under a far-field setting; i.e., the user and the anchors are in the far field of each other. For example, Shahmansoori et al. [11] proposed a three-stage localization algorithm that improves the simultaneous orthogonal matching pursuit algorithm and the space alternating generalized expectation-maximization algorithm, and exploited the sparsity of millimeter-wave multiple-input multiple-output (MIMO) channels to simplify the channel model. Hu et al. [12] presented a two-dimensional (2D) localization algorithm based on the celebrated estimating signal parameters via rotational invariance technique (ESPRIT), which estimates user positions by constructing signal subspaces.
@@
Some studies [13, 14] have explored the impact of realistic hardware impairments, including the CFO and PN, with no RIS considered. For instance, Meles et al. [13] simulated the effects of CFO and PN on angle-of-arrival (AOA) estimation. Chen et al. [14] considered the impacts of CFO, PN, mutual coupling, and power amplifier nonlinearity on 6G localization, and performed analysis using the derived misspecified Cramér-Rao lower bound (CRLB). However, neither of them provided an estimation algorithm to handle the hardware impairments.
@@
Some other works [15, 16, 17, 18] have delved into localization algorithm design for scenarios with either CFO or PN. On the one hand, the CFO was studied in [15, 16]. Qiu et al. [15] designed an maximum likelihood estimation (MLE)-based weighted iterative algorithm for the joint estimation of CFO and AOA, yet the coupling of CFO and position parameters in each iteration incurs high complexity. Zakariyya et al. [16] proposed a high-resolution multiple signal classification (MUSIC) algorithm for localization, but did not utilize the CFO estimates in localization, leading to a loss in localization accuracy. On the other hand, the PN was considered in [17, 18]. Xiao et al. [17] combined the expectation maximization (EM) and MUSIC algorithms for joint estimation of AOA and PN, but only considered a single-carrier system under a 2D setting. Zheng et al. [18] formulated a maximum a-posteriori (MAP) problem iteratively solved using the block majorization minimization (MM) algorithm after approximation, where the PN was simplified as a Gaussian variable. However, the above methods have not considered the use of RISs nor addressed the signal blockage by obstacles.
RIS has been integrated into localization systems to address the problem of signal blockage, but little consideration has been given to the above-mentioned hardware impairments, i.e., CFO and PN. From the perspective of user mobility, existing studies can be divided between static localization [19, 20, 21] and dynamic localization [22, 23]. The former encompasses scenarios where the users undergo minimal movement within a coherent time. The latter involves scenarios when the users move at a relatively high speed, and the Doppler shift cannot be overlooked. From the perspective of localization distance, existing research can be divided between far-field localization [21, 23, 24] and near-field localization [20, 21, 23, 25]. In the far-field case, the user-RIS distance exceeds Fraunhofer distance [26], and the signal can be approximated as a planar wave. Conversely, in the near-field case, the signal cannot be treated as parallel propagation and should be considered as a spherical wave.
@@
For example, Luan et al. [27] focused on the near-field static 2D localization scenario, and proposed a robust optimization algorithm to configure the RIS phase shifts for improving localization accuracy. The transmit powers were also optimized in the considered localization system. Ozturk et al. [28] further introduced a more practical RIS model with phase-dependent amplitude variations in the near-field circumstance. The misspecified CRLB was derived and an approximated mismatched maximum likelihood-based algorithm was proposed for joint UE localization and RIS parameter estimation. The system model in this paper also belongs to the near-field static category. Although we adopt the classical ideal RIS phase shift model with unit amplitude for simplicity, we first take the CFO and PN into consideration. Elzanaty et al. [23] studied RIS-assisted localization for multi-antenna mobile users. They proposed a three-dimensional (3D) model that applies to both far-field and near-field scenarios, and analyzed the position and orientation error bounds. Lin et al. [29] introduced a novel twin-RIS configuration comprising two RISs with relative spatial rotations. To extract the parameters of the cascaded channel, the authors employed array signal processing and atomic norm denoising techniques. They established nonlinear equations exploiting the characteristics of the twin-RIS structure to recover the multipath parameters. User localization for static users in the far-field scenario was then performed based on the estimated channel parameters. Teng et al. [22] introduced a new message-passing algorithm called BULT to locate mobile users in the far-field scenario. The Bayesian CRLB was derived for BULT.
@@
The existing RIS-assisted localization studies have assumed ideal channel conditions without considering realistic hardware impairments. Only several studies have considered modeling some specific impairments. For example, the studies in [23, 24, 30] took into account clock offset, and the work in [30] considered phase offsets. Zhang et al. [31] considered the CFO while overlooking the effects of PN. They proposed a MUSIC-based algorithm that directly estimates the position parameters together with the CFO. However, the algorithm is only suitable for far-field scenarios. To the best of our knowledge, no existing research has specifically analyzed the impact of both CFO and PN on RIS-assisted near-field localization systems.
@@ I-C Contributions
In this paper, we investigate a RIS-assisted OFDM system, where practical CFO and PN convolve with the data symbols, causing signal constellation rotation and ICI and, consequently, hindering user localization in near-field environments. The key contributions of this paper are outlined as follows:
  • We investigate a problem of static near-field localization in a RIS-aided OFDM system, while taking into account the CFO and PN. We propose an alternating optimization (AO)-based iterative estimation approach, where the MAP criterion and gradient descent (GD) algorithm are employed with closed-form expressions derived for the CFO and PN estimation.
@@
Figure 1: The considered RIS-aided 3D near-field localization system, including an AN, a UE, and an $N_{R}$-element RIS. The LoS link between the AN and UE is obstructed. The UE utilizes the signals transmitted by the AN and reflected by the RIS to localize its position.
@@
$N_{R}$ Number of RIS elements
@@
$f_{c}$ Carrier frequency of the system
@@
$U$ Position sampling amount
@@
$\bm{p}_{a}$ Coordinate of the AN
$\bm{p}_{r}$ Coordinate of the reference point of the RIS
$\bm{p}_{u}$ Coordinate of the UE position
@@ II System Model
The considered RIS-assisted OFDM downlink wireless localization system is composed of an anchor node (AN), a user equipment (UE), and an $N_{R}$-element RIS, as illustrated in Fig. 1. We make the assumption that the LoS path between the AN and UE is obstructed, and the channel state information (CSI) remains unchanged during localization, as in [21, 27, 30]. The positions of the AN and each RIS element, denoted as $\bm{p}_{a}$ and $\bm{p}_{r}=[x_{r},y_{r},z_{r}]^{\top}$, respectively, are known, where $r$ is the index of a RIS element. The position of UE, denoted as $\bm{p}_{u}=[x_{u},y_{u},z_{u}]^{\top}$, is to be estimated.
@@
The RIS elements are organized in a uniform planar array (UPA). We take the bottom-left corner of the RIS as the reference point and place it at the coordinate origin. The RIS is placed on the $yz$-plane, and both edges on the coordinate axes are located in the positive half-axis. Hence, the coordinates of the $r$-th RIS element can be represented as $\bm{p}_{r}=[0,d\times((r-1)\mod N_{R}),d\times\lceil\frac{r}{\sqrt{N_{R}}}-1\rceil]^{\top}$, where $d$ stands for the element spacing of the RIS, and is typically half of a wavelength. To satisfy the conditions of the near-field setup, both the AN-RIS distance, as well as the RIS-UE distance, should lie within the Fresnel region [26], i.e., the interval $[0.62\sqrt{D^{3}/\lambda},2D^{2}/\lambda]$, where $D=\sqrt{2}d(\sqrt{N_{R}}-1)$ represents the aperture of the RIS and $\lambda$ denotes the wavelength.
@@
Additionally, we adopt a polar coordinate system for estimating the user’s position, i.e., using $d_{Ru}$, $\phi_{az}$ and $\phi_{el}$, to express the user’s position, indicating the distance, azimuth angle, and elevation angle of UE relative to the coordinate origin, respectively. The relationships between these three parameters and $x_{u}$, $y_{u}$ and $z_{u}$ are given by
$d_{Ru}$ $=\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}};$ (1a)
$\phi_{az}$ $=\arctan\frac{y_{u}}{x_{u}};$ (1b)
$\phi_{el}$ $=\arccos\frac{z_{u}}{\sqrt{x_{u}^{2}+y_{u}^{2}+z_{u}^{2}}},$ (1c)
and, in turn, $x_{u}$, $y_{u}$ and $z_{u}$ can be written as
$x_{u}$ $=d_{Ru}\sin\phi_{el}\cos\phi_{az};$ (2a)
$y_{u}$ $=d_{Ru}\sin\phi_{el}\sin\phi_{az};$ (2b)
$z_{u}$ $=d_{Ru}\cos\phi_{el}.$ (2c)
@@
An OFDM symbol with a center carrier frequency of $f_{c}$, a signal bandwidth of $B$ and $N$ subcarriers is considered in this paper. Upon sampling at a rate of $T_{s}=\frac{1}{NB}$ and removing the cyclic prefix (CP) at the UE, the received time-domain signal $\bm{y}$ can be written as
@@
$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}_{LoS}+\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}_{NLoS}+\tilde{\bm{v}},$ (3)
@@
  • $\bm{h}_{LoS}$ and $\bm{h}_{NLoS}$ represent the LoS and non-LoS parts of the frequency-domain channel, respectively;
  • $\tilde{\bm{v}}$ is the additive Gaussian noise at the UE, and follows the distribution $\tilde{\bm{v}}\sim\mathcal{CN}(\bm{0}_{N\times 1},\sigma_{AWGN}^{2}\mathbf{I}_{N})$.
Assume that $\bm{h}_{NLoS}$ follows a known Gaussian distribution $\bm{h}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{I}_{N})$. Thus, $\bm{\mu}_{NLoS}\triangleq\mathbf{X}\bm{h}_{NLoS}$ also follows a known Gaussian distribution $\bm{\mu}_{NLoS}\sim\mathcal{CN}(0,\sigma_{NLoS}^{2}\mathbf{X}\mathbf{X}^{\rm H})$ with $\mathbf{X}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}$. Combine $\bm{\mu}_{NLoS}$ with the additive Gaussian noise $\tilde{\bm{v}}$. For conciseness, denote $\bm{h}_{LoS}$ with $\bm{h}$. The received signal is rewritten as
@@
where $\bm{v}\triangleq\bm{\mu}_{NLoS}+\tilde{\bm{v}}$ is Gaussian distributed, i.e, $\bm{v}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{N})$, with $\sigma^{2}=\sigma_{AWGN}^{2}+\sigma_{NLoS}^{2}$.
Each element of $\bm{h}$ can be written as $\bm{h}[n]=\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$, where $\sqrt{\rho_{aR,n,r}}=\frac{\sqrt{G_{a}}\lambda_{n}}{4\pi d_{ar}}$ and $\sqrt{\rho_{Ru,n,r}}=\frac{\sqrt{G_{t}}\lambda_{n}}{4\pi d_{ru}}$ are the free-space path losses from the AN to the $r$-th RIS element, and from the $r$-th RIS element to the UE, respectively. $\bm{w}[r]$ is the phase shift of the $r$-th RIS element, $\tau_{ar}=\frac{d_{ar}}{c}=\frac{\|\bm{p}_{r}-\bm{p}_{a}\|}{c}$ is the delay from the AN to the $r$-th RIS element, while $\tau_{ru}=\frac{d_{ru}}{c}=\frac{\|\bm{p}_{u}-\bm{p}_{r}\|}{c}$ denotes the delay from the $r$-th RIS element to the UE, with the light speed $c$.

### v4

@@ Abstract
Reconfigurable intelligent surface (RIS)-assisted communication systems have been extensively studied for providing high-precision location services. However, most studies have overlooked the impact of carrier frequency offset (CFO) and phase noise (PN) resulting from hardware impairments on localization. This paper presents a novel, alternating optimization (AO)-based algorithm to jointly estimate the CFO, PN, and user equipment (UE) position in orthogonal frequency division multiplexing (OFDM) systems, where, provided the UE position, closed-form expressions for the CFO and PN are derived per iteration, significantly reducing the complexity and enhancing the stability of the algorithm. Another important aspect is a new RIS phase shift optimization algorithm developed to minimize the analytical lower bound of localization accuracy, hence benefiting localization. The semidefinite relaxation method and Schur complement are utilized to convexify this challenging non-convex optimization problem to a semidefinite program. Simulations demonstrate the effectiveness of the proposed algorithms, with the localization accuracy enhanced by two orders of magnitude. The localization accuracy of the proposed algorithm is close to the analytical lower bound, with a root mean square error of lower than $\rm 10^{-2}\>m$.
@@ I-A Motivation and Challenges
With the development of the industrial Internet of Things, there is an increasing demand for centimeter-level localization accuracy in the sixth-generation (6G) mobile networks [1]. While global navigation satellite systems (GNSSs) have been widely applied to outdoor scenarios, they undergo insufficient localization accuracy and signal instability arising from obstacles [2, 3]. Targeting at areas with little access to satellite signals, wireless localization [4], also known as communication-assisted localization [5], has been widely studied. It can be potentially an integral part of integrated communication and sensing (ISAC) systems [6, 7] by utilizing communication devices and signals to achieve localization functionality.
While wireless localization systems have been extensively studied, few studies have considered the adverse effects of hardware impairments, such as carrier frequency offset (CFO) and phase noise (PN), on localization accuracy. CFO is a constant parameter due to the frequency offset between the transmitter and the receiver. Doppler shift can also lead to CFO. PN is a random process resulting from the fluctuation of the oscillators at the transmitters and receivers [8]. CFO and PN can result in signal constellation rotations in single carrier communication systems, or inter-carrier interference (ICI) in broadband communication systems, for example, orthogonal frequency division multiplexing (OFDM) systems. The overall communication performance and the localization accuracy can degrade significantly, due to the CFO and PN.
Another challenge arises from the increasingly likely blockages of signals, e.g., millimeter wave (mmWave) signals with the quasi-optical property, in many dense urban areas [9]. Reconfigurable intelligent surface (RIS) has been considered in communication systems to deal with the problem of signal blockage [10, 11]. Localization can be implemented with the help of virtual line-of-sight (LoS) paths provided by RISs even when the LoS paths are blocked, and can be improved via passive beamforming. In this sense, RIS-assisted user equipment (UE) localization is of practical interest, and could be particularly challenging with the further consideration of the CFO and PN. However, this has not been studied in the existing literature.
@@ I-B1 RIS-free localization with ideal hardware
Communication systems have been widely employed to assist in high-precision UE localization, typically under a far-field setting; i.e., the UE and the anchors are in the far field of each other. For example, Shahmansoori et al. [12] proposed a three-stage localization algorithm that improves the simultaneous orthogonal matching pursuit algorithm and the space alternating generalized expectation-maximization algorithm, and exploited the sparsity of millimeter-wave multiple-input multiple-output (MIMO) channels to simplify the channel model. Hu et al. [13] presented a two-dimensional (2D) localization algorithm based on the celebrated estimating signal parameters via rotational invariance technique (ESPRIT), which estimates UE positions by constructing signal subspaces.
I-B2 RIS-free localization under CFO and PN
@@
Some studies [14, 15] have explored the impact of realistic hardware impairments, including the CFO and PN, with no RIS considered. For instance, Meles et al. [14] simulated the effects of CFO and PN on angle-of-arrival (AOA) estimation. Nevertheless, CFO and PN were ignored when designing the algorithms and were only included in simulations, showing that the CFO and PN can significantly and adversely affect localization and tracking performance. Chen et al. [15] considered the impacts of CFO, PN, mutual coupling, and power amplifier nonlinearity on 6G localization, and derived the misspecified Cramér-Rao lower bound (CRLB). It was analytically concluded that the PN affects both angle and delay estimation, whereas the CFO has a more significant effect on angle estimation. However, neither of these works, i.e., [14, 15] have provided an estimation algorithm to handle hardware impairments.
@@
Some other works [16, 17, 18, 19] have delved into localization algorithm design with either CFO or PN. On the one hand, Qiu et al. [16] designed a maximum likelihood estimation (MLE)-based weighted iterative algorithm for the joint estimation of CFO and AOA. Zakariyya et al. [17] proposed a high-resolution enhancement of the multiple signal classification (MUSIC) algorithm for localization. While the algorithm balanced spectrum resolution and accuracy, the CFO was not estimated in localization, leading to a loss in localization accuracy. On the other hand, Xiao et al. [18] combined the expectation maximization (EM) and MUSIC algorithms for joint estimation of AOA and PN. At a typical signal-to-noise ratio (SNR) level, this algorithm can reduce the estimation error by over $55\%$ compared to its benchmarks. However, only a simple single-carrier system under a 2D far-field setting was considered. Zheng et al. [19] formulated a maximum a-posteriori (MAP) problem solved iteratively using the block majorization minimization (MM) algorithm after approximation. Under a medium or small PN, the algorithm improves localization and channel estimation. However, none of these methods [16, 17, 18, 19] have considered the use of RISs or addressed the signal blockage by obstacles.
I-B3 RIS-assisted localization with ideal hardware
@@
RIS has been integrated into localization systems to address the problem of signal blockage, but little consideration has been given to the above-mentioned hardware impairments, i.e., CFO and PN. From the perspective of localization distance, existing research can be divided between far-field localization [20, 21, 22] and near-field localization [23, 20, 21, 24]. In the far-field case, the UE-RIS distance exceeds Fraunhofer distance [25], and the signal can be approximated as a planar wave. Conversely, in the near-field case, the signal cannot be treated as parallel propagation and should be considered as a spherical wave. For example, Luan et al. [26] focused on the near-field 2D localization scenario, and proposed a robust optimization algorithm to configure the RIS phase shifts for improving localization accuracy. The transmit powers were also optimized in the considered localization system. Ozturk et al. [27] introduced a more practical RIS model with phase-dependent amplitude variations in the near-field circumstance. The mis-specified CRLB was derived and an approximated mismatched maximum likelihood-based algorithm was proposed for joint UE localization and RIS parameter estimation. Lin et al. [28] introduced a novel twin-RIS configuration comprising two RISs with relative spatial rotations. To extract the parameters of the cascaded channel, the authors employed array signal processing and atomic norm denoising techniques. They established nonlinear equations exploiting the characteristics of the twin-RIS structure to recover the multipath parameters. Localization for UEs in the far-field scenario was then performed based on the estimated channel parameters.
I-B4 RIS-assisted localization under CFO and PN
@@
Most existing RIS-assisted localization studies have assumed ideal channel conditions without considering realistic hardware impairments. Only several studies have considered modeling some specific impairments. For example, the studies in [21, 22, 29] took into account clock offset, and the work in [29] considered phase offsets. Zhang et al. [30] considered the CFO while overlooking the effects of PN. They proposed a MUSIC-based algorithm that directly estimates the position parameters together with the CFO. However, the algorithm is only suitable for far-field scenarios.
Unlike these existing works, we not only analytically quantify the impact of CFO and PN on localization accuracy, but also design efficient localization and RIS configuration algorithms. As a result, the localization accuracy can be improved by two orders of magnitude, compared to the existing methods. To the best of our knowledge, no existing research has specifically analyzed the impact of both CFO and PN on RIS-assisted near-field localization systems nor developed corresponding localization and RIS configuration algorithms.
@@ I-C Contributions
In this paper, we investigate a RIS-assisted OFDM system, where practical CFO and PN convolve with the data symbols, causing signal constellation rotation and ICI and, consequently, hindering UE localization in near-field environments. The key contributions of this paper are outlined as follows:
  • We investigate a problem of near-field localization in a RIS-aided OFDM system, while taking into account the CFO and PN. We propose an alternating optimization (AO)-based iterative estimation approach, where the MAP criterion and gradient descent (GD) algorithm are employed with closed-form expressions derived for the CFO and PN estimation.
@@
Figure 1: The considered RIS-aided 3D near-field localization system, including an AN, a UE, and an ${\color[rgb]{0,0,0}N_{\text{R}}}$-element RIS. The LoS link between the AN and UE is obstructed. The UE utilizes the signals transmitted by the AN and reflected by the RIS to localize its position.
@@
${\color[rgb]{0,0,0}N_{\text{R}}}$ Number of RIS elements
@@
${\color[rgb]{0,0,0}f_{\text{c}}}$ Carrier frequency of the system
@@
$\color[rgb]{0,0,0}S$ Position sampling amount
@@
${\color[rgb]{0,0,0}\bm{p}_{\text{a}}}$ Coordinate of the AN
$\bm{p}_{r}$ Coordinate of the $r$-th element of the RIS
$\bm{p}_{\color[rgb]{0,0,0}\text{u}}$ Coordinate of the UE position
@@ II System Model
The considered RIS-assisted OFDM downlink wireless localization system is composed of an anchor node (AN), a UE, and an $N_{\text{R}}$-element RIS, as illustrated in Fig. 1. We make the assumption that the LoS path between the AN and UE is obstructed, and the channel state information (CSI) remains unchanged during localization, as in [20, 26, 29]. Since both the AN and RIS are part of the network infrastructure, the positions of the AN and each RIS element, denoted as ${\color[rgb]{0,0,0}\bm{p}_{\text{a}}}$ and $\bm{p}_{r}=[x_{r},y_{r},z_{r}]^{\top}$, respectively, are known, as typically considered in RIS-assisted localization studies, e.g., [21, 26, 29, 31]. Here, $r$ is the index of a RIS element. The position of UE, denoted as $\bm{p}_{\color[rgb]{0,0,0}\text{u}}=[x_{\color[rgb]{0,0,0}\text{u}},y_{\color[rgb]{0,0,0}\text{u}},z_{\color[rgb]{0,0,0}\text{u}}]^{\top}$, is to be estimated.
@@
The RIS elements are organized in a uniform planar array (UPA). We take the bottom-left corner of the RIS as the reference point and place it at the coordinate origin. The RIS is placed on the $yz$-plane, and both edges on the coordinate axes are located in the positive half-axis. Hence, the coordinates of the $r$-th RIS element can be represented as $\bm{p}_{r}=[0,d\times(r\mod\sqrt{N_{\text{R}}}),d\times\lceil\frac{r+1}{\sqrt{N_{\text{R}}}}-1\rceil]^{\top}$, where $d$ stands for the element spacing of the RIS. We set $d$ to half-wavelength [23], [24], [33], because signals at different RIS elements may couple if $d$ is too small, or the RIS array’s sidelobe can be significant, leading to poor directionality for the RIS if $d$ is too large. To satisfy the conditions of the near-field setup, both the AN-RIS distance, as well as the RIS-UE distance, should lie within the Fresnel region [25], i.e., the interval $[0.62\sqrt{D^{3}/\lambda},2D^{2}/\lambda]$, where $D=\sqrt{2}d(\sqrt{{\color[rgb]{0,0,0}N_{\text{R}}}}-1)$ represents the aperture of the RIS and $\lambda$ denotes the wavelength.
@@
Additionally, we adopt a polar coordinate system for estimating the UE’s position, i.e., using ${\color[rgb]{0,0,0}d_{\text{R}}}$, ${\color[rgb]{0,0,0}\phi_{\text{az}}}$ and ${\color[rgb]{0,0,0}\phi_{\text{el}}}$, to express the UE’s position, indicating the distance, azimuth angle, and elevation angle of UE relative to the coordinate origin, respectively. The relationships between these three parameters and $x_{\color[rgb]{0,0,0}\text{u}}$, $y_{\color[rgb]{0,0,0}\text{u}}$ and $z_{\color[rgb]{0,0,0}\text{u}}$ are given by
${\color[rgb]{0,0,0}d_{\text{R}}}$ $=\sqrt{x_{\color[rgb]{0,0,0}\text{u}}^{2}+y_{\color[rgb]{0,0,0}\text{u}}^{2}+z_{\color[rgb]{0,0,0}\text{u}}^{2}};$ (1a)
${\color[rgb]{0,0,0}\phi_{\text{az}}}$ $=\arctan\frac{y_{\color[rgb]{0,0,0}\text{u}}}{x_{\color[rgb]{0,0,0}\text{u}}};$ (1b)
${\color[rgb]{0,0,0}\phi_{\text{el}}}$ $=\arccos\frac{z_{\color[rgb]{0,0,0}\text{u}}}{\sqrt{x_{\color[rgb]{0,0,0}\text{u}}^{2}+y_{\color[rgb]{0,0,0}\text{u}}^{2}+z_{\color[rgb]{0,0,0}\text{u}}^{2}}},$ (1c)
and, in turn, $x_{\color[rgb]{0,0,0}\text{u}}$, $y_{\color[rgb]{0,0,0}\text{u}}$ and $z_{\color[rgb]{0,0,0}\text{u}}$ can be written as
$x_{\color[rgb]{0,0,0}\text{u}}$ $={\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\cos{\color[rgb]{0,0,0}\phi_{\text{az}}};$ (2a)
$y_{\color[rgb]{0,0,0}\text{u}}$ $={\color[rgb]{0,0,0}d_{\text{R}}}\sin{\color[rgb]{0,0,0}\phi_{\text{el}}}\sin{\color[rgb]{0,0,0}\phi_{\text{az}}};$ (2b)
$z_{\color[rgb]{0,0,0}\text{u}}$ $={\color[rgb]{0,0,0}d_{\text{R}}}\cos{\color[rgb]{0,0,0}\phi_{\text{el}}}.$ (2c)
@@
An OFDM symbol with a center carrier frequency of ${\color[rgb]{0,0,0}f_{\text{c}}}$, a signal bandwidth of $B$ and $N$ subcarriers is considered in this paper. Upon sampling at a rate of ${\color[rgb]{0,0,0}T_{\text{s}}}=\frac{1}{NB}$ and removing the cyclic prefix (CP) at the UE, the received time-domain signal $\bm{y}$ can be written as
@@
$\bm{y}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}{\color[rgb]{0,0,0}\bm{h}_{\text{LoS}}}+\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}{\color[rgb]{0,0,0}\bm{h}_{\text{NLoS}}}+\tilde{\bm{v}},$ (3)
@@
  • ${\color[rgb]{0,0,0}\bm{h}_{\text{LoS}}}$ and ${\color[rgb]{0,0,0}\bm{h}_{\text{NLoS}}}$ represent the LoS and non-LoS parts of the frequency-domain channel, respectively [29, 34];
  • $\tilde{\bm{v}}$ is the additive Gaussian noise at the UE, and follows the distribution $\tilde{\bm{v}}\sim\mathcal{CN}(\bm{0}_{N\times 1},{\color[rgb]{0,0,0}\sigma_{\text{AWGN}}^{2}}\mathbf{I}_{N})$.
Assume that ${\color[rgb]{0,0,0}\bm{h}_{\text{NLoS}}}$ follows a known Gaussian distribution [35], i.e., ${\color[rgb]{0,0,0}\bm{h}_{\text{NLoS}}}\sim\mathcal{CN}({\color[rgb]{0,0,0}\bm{0}_{N\times 1}},{\color[rgb]{0,0,0}\sigma_{\text{NLoS}}^{2}}\mathbf{I}_{N})$. Thus, ${\color[rgb]{0,0,0}\bm{\mu}_{\text{NLoS}}}\triangleq\mathbf{X}{\color[rgb]{0,0,0}\bm{h}_{\text{NLoS}}}$ also follows a known Gaussian distribution ${\color[rgb]{0,0,0}\bm{\mu}_{\text{NLoS}}}\sim\mathcal{CN}({\color[rgb]{0,0,0}\bm{0}_{N\times 1}},{\color[rgb]{0,0,0}\sigma_{\text{NLoS}}^{2}}\mathbf{X}\mathbf{X}^{\rm H})$ with $\mathbf{X}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}$ and $\mathbf{X}\mathbf{X}^{\rm H}=NP\mathbf{I}_{N}$. Combine ${\color[rgb]{0,0,0}\bm{\mu}_{\text{NLoS}}}$ with the additive Gaussian noise $\tilde{\bm{v}}$. For conciseness, we rewrite ${\color[rgb]{0,0,0}\bm{h}_{\text{LoS}}}$ with $\bm{h}$ in the rest of this paper. The received signal is rewritten as
@@
where $\bm{v}\triangleq{\color[rgb]{0,0,0}\bm{\mu}_{\text{NLoS}}}+\tilde{\bm{v}}$ yields a Gaussian distribution, i.e, $\bm{v}\sim\mathcal{CN}({\color[rgb]{0,0,0}\bm{0}_{N\times 1}},\sigma^{2}\mathbf{I}_{N})$, with $\sigma^{2}={\color[rgb]{0,0,0}\sigma_{\text{AWGN}}^{2}}+{\color[rgb]{0,0,0}\sigma_{\text{NLoS}}^{2}NP}$.
Each element of $\bm{h}$ can be written as [26]
$\bm{h}[n]=\sum_{r=0}^{N_{\text{R}}-1}\sqrt{\rho^{n,r}_{\text{aR}}}\sqrt{\rho_{{\text{Ru}}}^{n,r}}\bm{w}[r]e^{-j2\pi\frac{nB}{N}(\tau_{\text{a}r}+{\color[rgb]{0,0,0}\tau_{r\text{u}}})},$ (7)
where ${\color[rgb]{0,0,0}\sqrt{\rho_{\text{aR}}^{n,r}}}=\frac{\sqrt{G_{\color[rgb]{0,0,0}\text{a}}}\lambda_{n}}{4\pi d_{{\color[rgb]{0,0,0}\text{a}}r}}$ and ${\color[rgb]{0,0,0}\sqrt{\rho_{\text{Ru}}^{n,r}}}=\frac{\sqrt{G_{\color[rgb]{0,0,0}\text{t}}}\lambda_{n}}{4\pi{\color[rgb]{0,0,0}d_{r\text{u}}}}$ are the free-space path losses from the AN to the $r$-th RIS element, and from the $r$-th RIS element to the UE, respectively; $\bm{w}[r]$ is the phase shift of the $r$-th RIS element; $\tau_{{\color[rgb]{0,0,0}\text{a}}r}=\frac{d_{{\color[rgb]{0,0,0}\text{a}}r}}{c}=\frac{\|\bm{p}_{r}-{\color[rgb]{0,0,0}\bm{p}_{\text{a}}}\|}{c}$ is the delay from the AN to the $r$-th RIS element, while ${\color[rgb]{0,0,0}\tau_{r\text{u}}}=\frac{{\color[rgb]{0,0,0}d_{r\text{u}}}}{c}=\frac{\|\bm{p}_{\color[rgb]{0,0,0}\text{u}}-\bm{p}_{r}\|}{c}$ is the delay from the $r$-th RIS element to the UE; $c$ is the speed of light.